# On The Collatz Conjecture: Topological and Ergodic Approach

### v2

@@ Abstract.
We study the Collatz function famously related to the Collatz Conjecture under the topological and ergodic perpectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the Alexandroff compactification of the established topology, we prove finiteness of cycles, which is a significant advance to the conjecture itself. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result for a general class of important maps.
@@ 1. Introduction
It has been extensively studied because of the famous Collatz Conjecture: For each $n\in\mathbb{N}$ there exists $k\in\mathbb{N}$ such that $f^{k}(n)=1$. In other words, every orbit enter the cycle $\{1,2,4\}$. This is the unique cycle known.
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles.
@@ 2. Main result
  • There exist a topology coarser than the discrete one and a $\sigma$-algebra with respect to which every recurrent point is periodic and the ergodic probabilities are the ones supported on the periodic orbits.
@@ 3.1. Invoking foundations
Let $\mathcal{T}_{\lambda}$ be a familiy of topologies on $\mathbb{N}$ and $\sigma(\mathcal{T}_{\lambda})$ their Borel $\sigma$-algebras. We have that
@@ 3.2. Trapping the orbits towards periodicity
The folllowing lemma build both a topology and its Borel $\sigma$-algebra making the Collatz map both measurable and predictable in the language of Ergodic Theory.
@@ 3.4. Peculiarity of the discrete topology
The unique topology making the Collatz function $f$ continuous and containing the colletion $\{\{n,2n\}\mid n\in\mathbb{N}\}$ is the discrete one.
@@
It is enough to show that if a topology satisfies the hypothesis then it contains all the singletons. In fact, we first show that it contains the even singletons $\{2k\}$. Once it contains $\{k,2k\}$ and $\{2k,4k\}$ it must contain the intersection $\{2k\}=\{k,2k\}\cap\{2k,4k\}$. Now, we show that it contains the odd singletons. Given an odd singleton $\{n\}$ we have that $n\in f^{-1}(\{(3n+1)/2,3n+1\})=\{n,3n+1\}$ which is open because by hypothsis $f$ is continuous and $\{(3n+1)/2,3n+1\}$ is open, since $n$ is odd and $3n+1$ is even. So, the following intersection is open $\{n\}=\{n,2n\}\cap\{n,3n+1\}.$ Once the topology contains all the singletons, it coincide with the discrete one. ∎
@@ 3.7. A key potential
The following lemma is key in the building of the bridge. It construct a special continuous and unbounded potentials that will be useful in the future.
@@
There exists a $\mathcal{T}$-continuous potential $\varphi:\mathbb{N}\to\mathbb{N}$ which is constant on periodic orbits whose integral with respect to an ergodic probability $\delta$ suppoted on $\mathcal{O}(x)$ is given by
@@ 3.8. The ergodic decomposition
This would be an consequence of the Ergodic Decomposition Theorem, but we cannot use it here because the topology is not even metrizable.
@@ 3.9. Finiteness of periodic orbits
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessesing at least one equilibrium state.
@@ 3.10. Uniqueness of periodic orbits: the final step
Uniqueness of periodic orbits is equivalent to every continuous and bounded potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessesing a unique equilibrium state.
@@
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bouned and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and the unique periodic orbit is $\{1,2,4\}$. ∎

### v3

@@ Abstract.
We study the Collatz function famously related to the Collatz Conjecture under the topological and ergodic perspectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the Alexandroff compactification of the established topology, we prove finiteness of cycles, which is a significant advance to the conjecture itself. We also present a proof of uniqueness of cycle and show that there is no divergent orbits. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result on the finiteness of orbits and no divergence for a general class of important maps.
@@ 1. Introduction
It has been extensively studied because of the famous Collatz Conjecture: For each $n\in\mathbb{N}$ there exists $k\in\mathbb{N}$ such that $f^{k}(n)=1$. In other words, every orbit enters the cycle $\{1,2,4\}$. This is the unique cycle known.
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles and later on, a proof of uniqueness of cycle. To prove the conjecture completely, we also prove that we cannot have divergent orbits. We also apply our technique for the Baker and Syracuse maps.
@@ 2. Main result
  • There exists a topology coarser than the discrete one and a $\sigma$-algebra with respect to which every recurrent point is periodic and the ergodic probabilities are the ones supported on the periodic orbits.
@@ 3.1. Invoking foundations
Let $\mathcal{T}_{\lambda}$ be a family of topologies on $\mathbb{N}$ and $\sigma(\mathcal{T}_{\lambda})$ their Borel $\sigma$-algebras. We have that
@@ 3.2. Trapping the orbits towards periodicity
The following lemma builds both a topology and its Borel $\sigma$-algebra making the Collatz map both measurable and predictable in the language of Ergodic Theory.
@@ 3.4. Peculiarity of the discrete topology
The unique topology making the Collatz function $f$ continuous and containing the collection $\{\{n,2n\}\mid n\in\mathbb{N}\}$ is the discrete one.
@@
It is enough to show that if a topology satisfies the hypothesis then it contains all the singletons. In fact, we first show that it contains the even singletons $\{2k\}$. Once it contains $\{k,2k\}$ and $\{2k,4k\}$ it must contain the intersection $\{2k\}=\{k,2k\}\cap\{2k,4k\}$. Now, we show that it contains the odd singletons. Given an odd singleton $\{n\}$ we have that $n\in f^{-1}(\{(3n+1)/2,3n+1\})=\{n,3n+1\}$ which is open because by hypothesis $f$ is continuous and $\{(3n+1)/2,3n+1\}$ is open, since $n$ is odd and $3n+1$ is even. So, the following intersection is open $\{n\}=\{n,2n\}\cap\{n,3n+1\}.$ Once the topology contains all the singletons, it coincides with the discrete one. ∎
@@ 3.7. A key potential
The following lemma is key in the building of the bridge. It construct a special continuous and unbounded potential that will be useful in the future.
@@
There exists a $\mathcal{T}$-continuous potential $\varphi:\mathbb{N}\to\mathbb{N}$ which is constant on periodic orbits whose integral with respect to an ergodic probability $\delta$ supported on $\mathcal{O}(x)$ is given by
@@ 3.8. The ergodic decomposition
This would be a consequence of the Ergodic Decomposition Theorem, but we cannot use it here because the topology is not even metrizable.
@@ 3.9. Finiteness of periodic orbits
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing at least one equilibrium state.
@@ 3.10. Uniqueness of periodic orbits: the final step
Uniqueness of periodic orbits is equivalent to every continuous and bounded potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing a unique equilibrium state.
@@
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bounded and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and the unique periodic orbit is $\{1,2,4\}$. ∎
@@ 4. Finiteness of Cycles
5. Uniqueness of Cycle
Theorem 17.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. The unique possible cycle is $\{1,2,4\}$.
Proof.
Let $\mathcal{O}_{1},\cdots,\mathcal{O}_{p}$ the finite list of cycles guaranteed by Theorem 16. Let us assume, by contradiction, then there exists $\mathcal{O}_{p}\neq\{1,2,4\}=\mathcal{O}_{1}$. As a consequence of the proof we know that the following set is clopen, compact and finite, with respect to the topology $\mathcal{T}$: $\mathcal{O}=\bigcup_{i=1}^{p}\mathcal{O}_{i}$.
Let $x:=\max\mathcal{O}$, which must be even because otherwise we would have $x<3x+1\in\mathcal{O}$. So $\{x,x/2\}\subset\mathcal{O}$. We have $x/2$ even because otherwise we would have $x<3(x/2)+1\in\mathcal{O}$. Also, there must exist an odd number $y\in\mathcal{O}$ such that $3y+1=x$ because otherwise $x<2x\in\mathcal{O}$. But we have the following geometric progression contained in $\mathcal{O}$: $\{y,2y,\cdots,2^{k}y\}$. But since $4y>3y+1=x=\max\mathcal{O}$, we must have $k=1$ and $\{y,2y\}\subset\mathcal{O}$. So, the following numbers are in the set $\mathcal{O}$:
$2y,y,x=3y+1,\frac{x}{2}=\frac{3y+1}{2}=y+\frac{y+1}{2}<y+y=2y.$
But we have
$y<\frac{3y+1}{2}<2y$
and the number $\frac{x}{2}=\frac{3y+1}{2}$ would be an even number in $\mathcal{O}$ between $y$ and $2y$. We also can find $z\in\mathcal{O}$ and odd number such that $3z+1=2y$. Also, the sequence $z,2z,\cdots,2^{m}z\in\mathcal{O}$. But since $8z>6z+2=2(3z+1)=2\cdot 2y=4y>x$ we must have $m=2$ and $z,2z,4z\in\mathcal{O}$. Moreover, we can find $w\in\mathcal{O}$ an odd number such that $3w+1=4z$ and $w,2w,\cdots,2^{q}w\in\mathcal{O}$. But $8w>6w+2=2(3w+1)=2\cdot 4z=8z>x$. So, we must have $q=2$ and $w,2w,4w\in\mathcal{O}$. We can find $t\in\mathcal{O}$ an odd number such that $3t+1=4w$. So $8t>6t+2=2(3t+1)=2\cdot 4w=8w>x$. By induction, we must find $a\in\mathcal{O}$ odd such that $a,2a,4a=x$. We have that
$z<y,z<w<t<\cdots<a.$
Since the cycles are pairwise disjoint, we can suppose that $\mathcal{O}_{p}$ is the unique cycle such that $x,y,z,w,t,a\in\mathcal{O}_{p}$. Moreover, the odd number $z=\min\mathcal{O}_{p}$. Then the orbit is
$\cdots\to\min\mathcal{O}_{p}=z\to 3z+1=2y\to y\to 3y+1=x=\max\mathcal{O}_{p}\to\cdots$
Then,
$x=3\frac{3z+1}{2}+1\implies x=\frac{9z+3}{2}+1\implies\max\mathcal{O}_{p}=\frac{9\min\mathcal{O}_{p}+3}{2}+1$
Also, we have $r$ geometric progressions with $3$ elements as $x=4a,2a,a$ and $1$ pair $2y,y$ and then the cycle $\mathcal{O}_{p}$ has $3r+2$ elements. We have an orbit as follows
$x=4a,2a,a,3a+1=4b,2b,b,3b+1=4c,2c,c.\cdots,4z,2z,z,3z+1=2y,y,3y+1=x.$
So, 4a4b=4a(3a+1)=a14a-4b=4a-(3a+1)=a-1, 4b4c=b1,4w4z=w14b-4c=b-1,4w-4z=w-1, which implies
$x=4a-4b+4b-4c+4c-\cdots+4w-4z+4z=$
$a-1+b-1+c-1+\cdots+w-1+z+3z=a+b+c+\cdots w+z-r+2y-1=$
$a+b+c+\cdots w+z+y-r+y-1=\frac{9z+3}{2}+1.$
The list $a,b,c,\cdots,w,z,y$ has $r+1$ odd numbers. If $r$ is odd, then $r+1$ is even and the sum $a+b+c+\cdots+w+z+y$ is even. So, yr1y-r-1 is odd and we get a contradiction because $x$ is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
Then, $r$ must be even and we have that the sum $a+b+c+\cdots+w+z+y$ is odd. But yr1y-r-1 is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
This contradiction shows that the cycle $\mathcal{O}_{p}$ cannot exist. We conclude that the unique cycle is $\mathcal{O}=\mathcal{O}_{1}=\{1,2,4\}$. ∎
@@ 6. Baker and Syracuse Maps
7. Divergent Orbits
Theorem 18.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. There is no divergent orbits.
Proof.
In fact, for every full orbit $\{f^{k}(n)\}_{k\in\mathbb{Z}}$, it is a clopen subset because it is a union of geometric progressions with ration $2$ and the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ is a union of finite geometric progressions with ration $2$. Its complement is also like this. We can take it as a covering. Since it is closed, it gets compact after the Alexandroff Compactification made in Section 4. Taking the union of geometric progressions itself as a covering, we can take a finite subcovering. So, the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ is a finite union of finite geometric progressions with ration $2$. It shows that the forward orbit is bounded and it must enter a cycle. Since the unique cycle is $\{1,2,4\}$ we must have every orbit reaching $1$. ∎
Remark 19.
We observe that the full orbit $\{f^{k}(n)\}_{k\in\mathbb{Z}}$ may contain and infinite geometric progression with ration $2$ but the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ cannot.